9-2m=19
7
I can't get the 7 under the 2m but it should read 2m over 7

Answer by Msgenes(10)

9 - 2m/7=19
9 - 2m/7-9 = 19 - 9
-2m/7 = 10
(-2m/7)/(-2/7) = 10/(-2/7)
When dividing fractions you multiply by the reciprocal
m = 10/1 * 7/2
m = 70/2
m = 35
Good Luck :)